# Temporal Cooperative Games

### v3

@@ Abstract
Classical cooperative game theory assumes that the worth of a coalition is determined solely by the set of agents involved. In practice, however, the worth may also depend on the order in which agents arrive. Motivated by such scenarios, we consider temporal cooperative games (TCG), where the worth of a set of agents depends on their order of arrival, i.e., the worth $v$ becomes a function of the sequence of agents $\pi$, rather than just the set $S$ of agents. This shift requires a fundamental rethinking of the desired axioms.
A key property in this temporal framework is the incentive for optimal arrival (I4OA), which encourages agents to join in the order that maximizes the total worth. Alongside this, we define two additional natural properties: online individual rationality (OIR), which incentivizes earlier agents to invite additional agents, and sequential efficiency (SE), which requires that the total worth for any sequence is fully distributed among its agents. We identify a class of reward-sharing mechanisms uniquely characterized by these three properties. The celebrated Shapley value does not directly apply here, and we construct natural analogs of the Shapley value in two variants: the sequential world, where rewards are defined for each sequence-player pair, and the extended world, where rewards are defined for each player alone. We show that properties efficiency, additivity, and null player uniquely determine the Shapley analogs in both worlds. Crucially, the Shapley analogs are disjoint from those satisfying the properties I4OA, OIR, and efficiency. The conflict persists even for the restricted classes of convex and simple temporal cooperative games.
Our results thus reveal a fundamental tension: when players arrive sequentially, reward-sharing mechanisms that satisfy desirable temporal properties must differ in nature from the Shapley analogs. This opens up a bigger research question of finding good solution concepts for TCGs.
@@ 1 Introduction
Organizations and institutions benefit from the complementary skill sets of their employees. However, the growth and overall value of an institution critically depend on the order in which employees join. Early joiners establish the vision and direction of the institution, which subsequently guides the hiring of core domain experts, followed by process enablers. Support staff and marketing personnel are then integrated in a sequence that aligns with the institution’s structure. A different order of joining can result in markedly different growth trajectories and institutional value. Classical cooperative games (Maschler et al., 2020) often overlook this phenomenon, as their worth functions are defined only for coalitions, ignoring the order of the agents within the coalition. While Nowak and Radzik (1994); Sánchez and Bergantiños (1997) incorporate agent order through generalized characteristic functions, their primary goal is to extend the Shapley value to that context. In contrast, we focus on cooperative games in which the worth is sequence-dependent, which we term temporal cooperative games (TCGs), and our primary objective is to study the properties desirable in this new space as well as their simultaneous satisfaction with Shapley-like properties.
The first two properties we define for TCGs are inspired by similar notions in Ge et al. (2024). They introduce incentive for early arrivals, which encourages players to join as soon as they arrive a notion suited for classical cooperative games with dynamic arrivals. In this paper, we aim to incentivize agents to join in the optimal order that maximizes institutional worth. We term this property incentive for optimal arrival (I4OA). Without it, agents might join in arbitrary order, driven, for example, by short-term salary gains for high-profile roles such as CEO or CTO, rather than by their contribution to institutional value, which is undesirable.
@@
The second property, online individual rationality (OIR), is identical to that in Ge et al. (2024). OIR ensures that agents joining earlier receive weakly better payoffs as subsequent agents join, thereby encouraging early agents to facilitate the arrival of additional participants.
@@
We also consider a standard third property, sequential efficiency (SE), which guarantees that the worth of any sequence is fully allocated among the agents in that sequence. This property is crucial when rewards must be distributed immediately upon agent arrivals, without waiting for future participants. In this paper, we characterize a class of solution concepts that satisfy these three properties.
@@
Ideally, one would like a solution concept that satisfies these properties while also adhering to Shapley-inspired axioms in TCGs. However, we prove that this is impossible: solution concepts that uniquely satisfy Shapley-inspired properties such as efficiency, additivity, and the null-player condition are disjoint from those that satisfy TCG-appropriate properties like I4OA, OIR, and efficiency (Figure˜6). Remarkably, this impossibility persists even for restricted TCG classes such as convex and simple games, highlighting a stark contrast with classical cooperative games.
@@ 1.1 Our Contributions
  • We study the framework of temporal cooperative games (TCGs), originally proposed by Nowak and Radzik (1994), where worth is assigned to sequences of agents rather than coalitions. We introduce properties that are natural in this setting, namely I4OA, OIR, and SE (Section˜2).
  • We design a class of reward-sharing mechanisms (SeqShare, Algorithm˜1) that uniquely satisfies these properties. Furthermore, we identify a condition, termed basis, that is necessary for their satisfaction (Lemma˜1).
  • To compare our results with Shapley’s axioms, we formulate suitable adaptations of additivity, null-player, and efficiency for TCGs, where solution concepts are defined for every sequence (Section˜4). We show that a Shapley-analogous solution concept, MargSol, uniquely satisfies these properties (Section˜5.1).
  • Since the Shapley value averages over sequences to give a single value to every player, we extend our solution concepts to define extended counterparts. We prove that extended additivity, null-player, and efficiency uniquely identify the Ext-Shap solution concept (Section˜5.2).
  • We prove that a property analogous to symmetry follows directly from the three earlier properties (Section˜6), underscoring both the richness of the domain and the strictness of the axioms in this setting.
  • Finally, we show that Ext-Shap is disjoint from the extended version of SeqShare, establishing that the desirable properties of TCGs are fundamentally in conflict with Shapley-inspired properties, even in special classes such as convex and simple games (Section˜7).
@@ 1.2 Related Work
Classical cooperative game theory originated in the pioneering work of Von Neumann and Morgenstern (1947). Soon after, foundational solution concepts such as the core (Gillies, 1959), the nucleolus (Schmeidler, 1969), and the Shapley value (Shapley, 1953) were introduced. The study of online coalition formation is relatively recent: Flammini et al. (2021) addressed the problem in its general form, Bullinger and Romen (2023) considered random arrivals, and Bullinger and Romen (2025) studied stability in this setting. In parallel, Lehrer and Scarsini (2013) examined dynamic cooperative games in which the worth of a coalition varies over time, while Habis and Jean-Jacques Herings (2010); Kranich et al. (2005) extended the concept of the core to dynamic environments. Some other works consider the characterisation of solutions concepts that satisfy modification of the Shapley properties for classical cooperative games. Van Den Brink (2002) use Efficiency Null Player and an additional property fairness defined appropriately for cooperative games, to characterize the Shapley value. Casajus and Huettner (2013) modify the null player property to obtain a class of solutions. Several other works propose alternative properties that characterize the Shapley Value (Hamiache, 2001; Casajus, 2014).
@@
Online cooperative games, where agents arrive sequentially and are incentivised to satisfy properties such as immediate participation, have been studied in monotone coalitional games by Ge et al. (2024) and in cost-sharing games by Zhang et al. (2025a). Axiomatic approaches for online cooperative games were explored by Aziz et al. (2025), while Zhang et al. (2025b) developed stable online coalition formation mechanisms aimed at maximizing social welfare. However, in all these models, the worth is still defined for coalitions of agents, not for sequences. By contrast, our work focuses on cooperative games in which the worth depends explicitly on sequences, and one of the objectives is to incentivize an optimal arrival sequence rather than merely incentivizing early arrival.
The idea of generalized characteristic functions defined over ordered subsets was first proposed by Nowak and Radzik (1994) and subsequently extended by Sánchez and Bergantiños (1997); Sánchez and Bergantiños (1999); Bergantiños and Sánchez (2001). While these functions define worth in a sequence-dependent manner, the associated solution concepts remain sequence-independent. In particular, Nowak and Radzik (1994) extended Shapley-like properties, such as additivity, null player, and efficiency, to temporal cooperative games (TCGs) and proposed a solution concept that uniquely satisfies them. These results are closely related to some of ours, which we obtained independently before becoming aware of their work. Our focus, however, is on identifying and characterizing properties that are desirable in TCGs and demonstrating their conflict with Shapley-inspired properties, even within restricted settings such as convex and simple games. Furthermore, we establish connections between sequence-dependent and sequence-independent solution concepts, leading to two unique solutions: MargSol and Ext-Shap. Finally, Sánchez and Bergantiños (1997) proposed a weaker version of the null-player property than ours (for details, see Section˜4) and required the introduction of symmetry to obtain a unique solution that satisfies all four properties, which we do not. So, even though this literature is closest to our work, there remain significant differences in both modeling choices and results.
@@ 2 Preliminaries
We call $\pi^{\prime}$ to be a prefix of $\pi$ if the first $|P(\pi^{\prime})|$ players of $\pi$ are the same as that of $\pi^{\prime}$, i.e., $\pi(i)=\pi^{\prime}(i),\forall i=1,2,\ldots,|P(\pi^{\prime})|$. A prefix $\pi^{\prime}$ of $\pi$ is denoted as $\pi^{\prime}\sqsubset\pi$. Note that $|P(\pi^{\prime})|<|P(\pi)|$, otherwise they become the same sequence. We assume that the characteristic function is monotone, i.e., $v(\pi^{\prime})\leqslant v(\pi)$ for all $\pi^{\prime}\sqsubset\pi$. We represent the predecessor of $i$ in $\pi$ with $\pi_{i}$ which denotes the longest prefix of $\pi$ not containing $i$. A sequence $\pi$ is a full sequence if it contains all the players, $P(\pi)=N$. Denote the set of all monotone characteristic functions by $V$.
The goal of this paper is to find desirable axioms under this setting and obtain a solution concept $\phi_{i},i\in N$, which divides the worth among the players and satisfies these axioms. The term $\phi_{i}(\pi,v)$ denotes the share of value to agent $i\in P(\pi)$ when arriving in the sequence $\pi$ in the TCG $\langle N,v\rangle$. We denote by $\phi(\pi,v)$ the vector $(\phi_{i}(\pi,v),i\in N)$, with the convention that the solution concept assigns zero reward to the agents outside $\pi$, i.e., $\phi_{j}(\pi,v)=0,\ \forall j\in N\setminus P(\pi)$. For the rest of the paper, we use $\phi(\pi)$ to denote the reward vector, dropping $v$ when it is clear from context. We denote the optimal sequence $\pi^{*}(v)$ to be the one that maximizes the worth of this game, i.e., $\pi^{*}(v)\in\operatorname*{\arg\!\max}_{\pi\in\Pi}v(\pi)$. WLOG, we assume that the tie-breaking rule is such that $P(\pi^{*}(v))=N$ (with arbitrary tie-breaking among the full length sequences), since the game is monotone. However, our results hold true even when $P(\pi^{*}(v))\subset N$, and the proof can be easily adapted.
@@
A solution concept $\phi$ is online individually rational (OIR) if for every TCG $\langle N,v\rangle$
$\phi_{i}(\pi,v)\leqslant\phi_{i}(\pi^{\prime},v),\forall i\in P(\pi),\forall\pi\sqsubset\pi^{\prime},\text{ where }\pi,\pi^{\prime}\in\Pi.$
The sequential nature of the game also creates the possibility of the players to stop arriving. Also the next player’s arrival is uncertain beforehand. Therefore, the solution concept should ensure the the worth of a given sequence $\pi$ is given out to all the players in $P(\pi)$ for every sequence $\pi$.
@@
A solution concept $\phi$ is sequentially efficient if for every TCG $\langle N,v\rangle$
$\sum_{i\in P(\pi)}\phi_{i}(\pi,v)=v(\pi),\forall\pi\in\Pi.$
Since the number of agents are finite, every TCG will have an optimal sequence that generates the most value. We denote the optimal sequence of a TCG $\langle N,v\rangle$ by $\pi^{*}(v)$. Our next property incentivizes players to arrive in this optimal order. We use $\pi^{*}$ to denote the optimal sequence when the game $\langle N,v\rangle$ is clear from the context.
@@
A solution concept $\phi$ satisfies incentive for optimal arrival (I4OA) if for every TCG $\langle N,v\rangle$
$\phi_{i}(\pi^{*}(v),v)\geqslant\phi_{i}(\pi,v),\forall\pi\in\Pi,i\in N.$
In the first part of this paper, we characterize reward sharing mechanisms that satisfy OIR, I4OA, and SE for TCGs. In the second part, we define properties analogous to the celebrated Shapley properties in classical cooperative games and characterize the solution concepts that satisfy them.

### v4

# Sequential Solution Concepts in Cooperative Games with Generalized Characteristic Functions

@@ Abstract
Motivated by the fact that the worth of a coalition may depend on the order in which agents arrive, Nowak and Radzik (1994) (NR) introduced cooperative games with generalized characteristic functions. We study such temporal cooperative games (TCGs), where the worth function $v$ is defined on sequences of agents $\pi$ rather than sets $S$. This order sensitivity necessitates a re-examination of axioms for reward sharing. NR and subsequent work proposed several axioms; the resulting solution concepts are still inherently order-oblivious and closely tied to the Shapley value.
In contrast, we focus on sequential solution concepts that explicitly depend on the realized order $\pi$. We study reward-sharing mechanisms satisfying incentive for optimal arrival (I4OA), which promotes orders maximizing total worth; online individual rationality (OIR), which ensures agents are not harmed by later arrivals; and sequential efficiency (SE), which requires that the worth of any sequence is fully distributed among its agents. These axioms are intrinsic to TCGs, and we characterize a class of reward-sharing mechanisms uniquely determined by them. The classical Shapley value does not directly extend to this setting. We therefore construct natural Shapley analogs in two worlds: a sequential world, where rewards are defined for each sequence–agent pair, and an extended world, where rewards are defined per agent, consistent with the NR framework. In both cases, the axioms of efficiency, additivity, and null player uniquely characterize the corresponding Shapley analogs. But, these Shapley analogs are disjoint from the class of solutions satisfying the sequential axioms, even for convex and simple TCGs.
Our results reveal a fundamental tension in temporal cooperative games: when order matters, solution concepts satisfying natural sequential incentives must differ structurally from Shapley-based allocations, motivating new axiomatic foundations for sequential reward sharing.
@@ 1 Introduction
Organizations and institutions benefit from the complementary skill sets of their employees. However, the growth and overall value of an institution critically depend on the order in which employees join. Early joiners establish the vision and direction of the institution, which subsequently guides the hiring of core domain experts, followed by process enablers. Support staff and marketing personnel are then integrated in a sequence that aligns with the institution’s structure. A different order of joining can result in markedly different growth trajectories and institutional value. Classical cooperative games (Maschler et al., 2020) often overlook this phenomenon, as their worth functions are defined only for coalitions, ignoring the order of the agents within the coalition. Early works by Nowak and Radzik (1994); Sánchez and Bergantiños (1997) incorporate agent order through generalized characteristic functions (GCF). A series of follow-up papers (Van den Brink et al., 2014; Feng et al., 2013; Yuan and Driessen, 2013; Li et al., 2024; Michalak et al., 2014; Zou et al., 2020) consider extending or modifying some of their axioms and either characterize new solution concepts or provide alternative proofs. However, a primary feature of all these papers is that they consider solution concepts that are order-oblivious. The axiomatic characterization results try to extend the Shapley value for a player that does not depend on the order of arrival. In contrast, we focus on solution concepts that are order-sensitive, and our primary objective is to study the desirable properties of the order-sensitive solution concepts as well as their simultaneous satisfaction with Shapley-like properties that are order-sensitive. We will be referring to the cooperative games with generalized characteristic functions as temporal cooperative games (TCG) to distinguish them from classical cooperative games (CG) in the rest of the paper.
TCGs necessitate a new solution concept along with a corresponding set of desirable axioms. We argue that such a solution concept must be order-sensitive, since the same agent appearing in different positions of a sequence can induce different worths in TCGs, and therefore should receive different rewards. Motivated by this observation, we introduce three fundamental sequential properties, inspired by related notions in Ge et al. (2024). The first property, incentive for optimal arrival (I4OA), incentivizes agents to join in the optimal order that maximizes institutional worth.
@@
The second property, online individual rationality (OIR), coincides with the notion introduced in Ge et al. (2024). OIR ensures that agents who join earlier receive weakly higher payoffs as additional agents arrive, thereby encouraging early participants to facilitate the arrival of others.
@@
We also consider a standard third property, sequential efficiency (SE), which guarantees that the worth of any sequence is fully allocated among the agents in that sequence. This property is particularly important in settings where rewards must be distributed immediately upon agent arrivals, without waiting for future participants. In this paper, we characterize a class of solution concepts that satisfy these three properties. We further provide an approach illustrating how additional properties may restrict this class toward a unique solution concept.
@@
Ideally, one would like such solution concepts to also satisfy Shapley-inspired order-sensitive axioms in TCGs. However, we prove that this is impossible: solution concepts that uniquely satisfy Shapley-inspired properties such as efficiency, additivity, and the null-player condition are disjoint from those satisfying TCG-appropriate properties such as I4OA, OIR, and efficiency (Figure˜3). Remarkably, this impossibility result persists even for restricted classes of TCGs, including convex and simple games, thereby underscoring a sharp departure from the classical cooperative game setting.
@@ 1.1 Our Contributions
  • We study cooperative games with generalized characteristic functions (GCF) introduced by Nowak and Radzik (1994), which we call temporal cooperative games (TCGs). We formalize axioms natural to this order-sensitive setting: incentive for optimal arrival (I4OA), online individual rationality (OIR), and sequential efficiency (SE) (Section˜2).
  • A key contribution is the introduction of sequential solution concepts (Section˜2). Unlike existing GCF literature, which yields order-oblivious reward shares to the players, our framework assigns rewards to every (sequence, player) pair, arguing that order-sensitive environments require order-sensitive solution concepts.
  • We characterize all solution concepts satisfying I4OA, OIR, and SE via a class of reward-sharing mechanisms SeqShare (Algorithm˜1), and identify a necessary condition, termed basis, for their satisfaction (Lemma˜1).
  • We formulate order-sensitive analogs of Shapley’s axioms for TCGs: efficiency, additivity, and null-player (Section˜4), and show that they uniquely characterize a Shapley-analogous sequential solution concept MargSol (Section˜5.1); symmetry follows as a consequence (Section˜6).
  • To connect with order-oblivious solutions, we define extended solution concepts by averaging over sequences. The extended axioms uniquely identify Ext-Shap (Section˜5.2), aligning exactly with the results in existing GCF literature.
  • Finally, we show that the extended solutions induced by SeqShare are disjoint from Ext-Shap, establishing a fundamental incompatibility between desirable sequential properties and Shapley-inspired axioms, even for convex and simple games (Section˜7).
@@ 1.2 Related Work
Classical cooperative game theory originated with the seminal work of Von Neumann and Morgenstern (1947). Soon thereafter, foundational solution concepts such as the core (Gillies, 1959), the nucleolus (Schmeidler, 1969), and the Shapley value (Shapley, 1953) were introduced. The study of online coalition formation is comparatively recent: Flammini et al. (2021) addressed the problem in its general form, Bullinger and Romen (2023) analyzed settings with random arrivals, and Bullinger and Romen (2025) examined stability in this context. In parallel, Lehrer and Scarsini (2013) studied dynamic cooperative games in which the worth of a coalition evolves over time, while Habis and Jean-Jacques Herings (2010); Kranich et al. (2005) extended the notion of the core to dynamic environments. Other strands of the literature focus on characterizing solution concepts that satisfy variants of the Shapley axioms for classical cooperative games. In particular, Van Den Brink (2002) employ efficiency, the null-player property, and fairness (appropriately defined for cooperative games) to characterize the Shapley value, while Casajus and Huettner (2013) modify the null-player property to obtain a broader class of solutions. Several other works propose alternative axiom systems characterizing the Shapley value (Hamiache, 2001; Casajus, 2014).
@@
Online cooperative games, in which agents arrive sequentially and are incentivized to satisfy properties such as immediate participation, have been studied for monotone coalitional games by Ge et al. (2024) and for cost-sharing games by Zhang et al. (2025a). Axiomatic frameworks for online cooperative games were developed by Aziz et al. (2025), while Zhang et al. (2025b) proposed stable online coalition formation mechanisms aimed at maximizing social welfare. However, across all these models, worth is still defined over coalitions of agents rather than over sequences.
The notion of generalized characteristic functions defined on ordered subsets was first introduced by Nowak and Radzik (1994) and subsequently extended by Sánchez and Bergantiños (1997, 1999); Bergantiños and Sánchez (2001). Although these functions allow coalition worth to depend on order, the resulting solution concepts are sequence-independent. We refer to these as the NR and SB values, respectively. In particular, Nowak and Radzik (1994) extended Shapley-style axioms—such as additivity, the null-player condition, and efficiency—to temporal cooperative games (TCGs) and proposed a solution concept that uniquely satisfies them. Sánchez and Bergantiños (1997) introduced a weaker version of the null-player property than the one we consider (see Section˜4 for details) and required the additional assumption of symmetry to obtain a unique solution satisfying all four properties, which we do not impose.
In subsequent work on generalized characteristic functions, Yuan and Driessen (2013) characterize the SB value as the outcome of a bargaining game, while Feng et al. (2013) provide a characterization based on three properties, namely continuity, associated consistency, and the inessential game property. Van den Brink et al. (2014) introduce the order monotonicity axiom and characterize a class of solutions that includes the SB and NR values as extreme cases. Zou et al. (2020) and Li et al. (2024) study settings with restrictions on the order of agent arrivals and characterize solutions for these constrained generalized games. Finally, Michalak et al. (2014) implement generalized games as perfect-information games that admit the NR and SB values as subgame-perfect Nash equilibria.
The fundamental distinction between these works and ours lies in our focus on order-sensitive solution concepts. This perspective leads to two key departures: (1) the axioms of interest differ and are motivated by settings in which order-sensitive solutions are natural, and (2) the required properties are more stringent in the Shapley framework, as they must hold for every sequence $\pi$. Consequently, our main results are also different: (1) a characterization of order-sensitive axioms tailored to TCGs (Theorem˜1), (2) a characterization of Shapley-like yet order-sensitive axioms for TCGs (Theorem˜2), and (3) a demonstration of the disjointness of these two classes (Section˜7).
@@ 2 Preliminaries
We call $\pi^{\prime}$ to be a prefix of $\pi$ if the first $|P(\pi^{\prime})|$ players of $\pi$ are the same as that of $\pi^{\prime}$, i.e., $\pi(i)=\pi^{\prime}(i),\forall i=1,2,\ldots,|P(\pi^{\prime})|$. A prefix $\pi^{\prime}$ of $\pi$ is denoted as $\pi^{\prime}\sqsubset\pi$. Note that $|P(\pi^{\prime})|<|P(\pi)|$, otherwise they become the same sequence. We assume that the characteristic function is monotone, i.e., $v(\pi^{\prime})\leqslant v(\pi)$ for all $\pi^{\prime}\sqsubset\pi$. We represent the predecessor of $i$ in $\pi$ with $\pi_{i}$, which denotes the longest prefix of $\pi$ not containing $i$. A sequence $\pi$ is a full sequence if it contains all the players, $P(\pi)=N$. Denote the set of all monotone characteristic functions by $V$.
The focus of this paper is to find desirable axioms under this setting and obtain an order-sensitive solution concept $\phi_{i},i\in N$, which divides the worth among the players for every sequence and satisfies these axioms. The term $\phi_{i}(\pi,v)$ denotes the share of value to agent $i\in P(\pi)$ when arriving in the sequence $\pi$ in the TCG $\langle N,v\rangle$. We denote by $\phi(\pi,v)$ the vector $(\phi_{i}(\pi,v),i\in N)$, with the convention that the solution concept assigns zero reward to the agents outside $\pi$, i.e., $\phi_{j}(\pi,v)=0,\ \forall j\in N\setminus P(\pi)$. For the rest of the paper, we use $\phi(\pi)$ to denote the reward vector, dropping $v$ when it is clear from context. We denote the optimal sequence $\pi^{*}(v)$ to be the one that maximizes the worth of this game, i.e., $\pi^{*}(v)\in\operatorname*{\arg\!\max}_{\pi\in\Pi}v(\pi)$. WLOG, we assume that the tie-breaking rule is such that $P(\pi^{*}(v))=N$ (with arbitrary tie-breaking among the full-length sequences), since the game is monotone. However, our results hold true even when $P(\pi^{*}(v))\subset N$, and the proof can be easily adapted.
@@
A solution concept $\phi$ is online individually rational (OIR) if for every TCG $\langle N,v\rangle$ $\phi_{i}(\pi,v)\leqslant\phi_{i}(\pi^{\prime},v),\forall i\in P(\pi),\forall\pi\sqsubset\pi^{\prime},\text{ where }\pi,\pi^{\prime}\in\Pi.$
The sequential nature of the game also creates the possibility of the players to stop arriving. Also, the next player’s arrival is uncertain beforehand. Therefore, the solution concept should ensure the worth of a given sequence $\pi$ is given out to all the players in $P(\pi)$ for every sequence $\pi$.
@@
A solution concept $\phi$ is sequentially efficient if for every TCG $\langle N,v\rangle$ $\sum_{i\in P(\pi)}\phi_{i}(\pi,v)=v(\pi),\forall\pi\in\Pi.$
Since the number of agents is finite, every TCG will have an optimal sequence that generates the most value. We denote the optimal sequence of a TCG $\langle N,v\rangle$ by $\pi^{*}(v)$. Our next property incentivizes players to arrive in this optimal order. We use $\pi^{*}$ to denote the optimal sequence when the game $\langle N,v\rangle$ is clear from the context.
@@
A solution concept $\phi$ satisfies incentive for optimal arrival (I4OA) if for every TCG $\langle N,v\rangle$ $\phi_{i}(\pi^{*}(v),v)\geqslant\phi_{i}(\pi,v),\forall\pi\in\Pi,i\in N.$
In the first part of this paper, we characterize reward-sharing mechanisms that satisfy OIR, I4OA, and SE for TCGs. In the second part, we define properties analogous to the celebrated Shapley properties in classical cooperative games and characterize the solution concepts that satisfy them.